# Noncritical Conformal Gravity and Four-Dimensional Liouville Theory

### v1

@@ 2 Conformal anomaly from the conformal gravity
where $\delta\lambda$ is an anticommuting parameter and $c$ is a ghost field for the conformal transformation. The BRST transformation of the ghost is determined by the requirement that the BRST transformation (2.2) be nilpotent. This gives
@@
where anticommuting property of the ghosts is used so that bc=cbbc=-cb. [^1] If we use the transformation (2.4), the same procedure gives $S_{\rm GF+FP}=\int d^{4}x\sqrt{g}\left[-\frac{\alpha}{2}\left(B+cb-\frac{1}{\alpha}R\right)^{2}+\frac{1}{2\alpha}R^{2}+3b\square c\right].$ So practically both transformations give the same result. The original conformal transformation is broken by the gauge fixing term, leaving only the global BRST invariance. Thus we may adopt transformations for the antighost and auxiliary fields different from those suggested by the conformal transformation as long as we consider the proper transformation for the other physical fields. It is interesting that there is no RbcRbc term in the action (2.9) even though $b$ and $c$ are (fermionic) scalar fields, in contrast to usual bosonic scalar fields (see Appendix B).
@@ 3 Lessons from two-dimensional quantum gravity for conformal gravity – Integrating the trace anomaly
$-\frac{c}{48\pi\epsilon_{2}}\int d^{D}x\sqrt{{\hat{g}}}\,e^{(D-2)\phi/2}\left(\hat{R}-(D-1)\hat{\square}\phi-\frac{(D-1)(D-2)}{4}(\partial_{\mu}\sigma)^{2}\right),$ (3.13)
@@ 4.1 Integrating trace anomaly
The Weyl tensor squared in $D$ diemnsions is
@@
$\sqrt{g}\,C_{\mu\nu\lambda\sigma}^{2}\to\sqrt{g}\,e^{((D-4)\phi/2}\hat{C}_{\mu\nu\lambda\rho}^{2}.$ (4.4)
@@
$\frac{1}{\epsilon}\int d^{D}x\sqrt{g}\,C_{\mu\nu\lambda\sigma}^{2},$ (4.5)
@@
$\int d^{D}x\sqrt{{\hat{g}}}\,\phi\,\hat{C}_{\mu\nu\lambda\sigma}^{2}.$ (4.6)
@@
$-\frac{1}{\epsilon}\int d^{D}x\sqrt{g}\,R_{GB}^{2},$ (4.10)
@@
$\int d^{4}x\Big[\phi\,\hat{R}_{\rm GB}^{2}-2\hat{G}^{\mu\nu}\hat{\nabla}_{\mu}\phi\hat{\nabla}_{\nu}\phi+2(\hat{\nabla}^{\mu}\hat{\nabla}^{\nu}\phi)\partial_{\mu}\phi\partial_{\nu}\phi-\frac{1}{4}\{(\partial_{\mu}\phi)^{2}\}^{2}\Big]$
$=\int d^{4}x\sqrt{g}\Big[\phi\,\hat{R}_{\rm GB}^{2}-2\hat{G}^{\mu\nu}\hat{\nabla}_{\mu}\phi\hat{\nabla}_{\nu}\phi-(\partial_{\mu}\phi)^{2}\hat{\square}\phi-\frac{1}{4}\{(\partial_{\mu}\phi)^{2}\}^{2}\Big].$ (4.11)
Collecting, our four-dimensional Liouville action whould be
$S_{\rm eff}\hskip-5.69054pt$ $=$ $\hskip-5.69054pt\int d^{4}x\sqrt{{\hat{g}}}\,\Big[\frac{c}{(4\pi)^{2}}\phi\,\hat{C}_{\mu\nu\lambda\sigma}^{2}-\frac{a}{(4\pi)^{2}}\Big(\phi\,\hat{R}_{\rm GB}^{2}-2\hat{G}^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi-(\hat{\square}\phi)(\partial_{\mu}\phi)^{2}-\frac{1}{4}\{(\partial_{\mu}\phi)^{2}\}^{2}\Big)\Big].$
@@ 4.2 Four-dimensional noncritical conformal gravity
$S_{4D,L}[{\hat{g}},\phi]\hskip-5.69054pt$ $=$ $\hskip-5.69054pt\int d^{4}x\sqrt{{\hat{g}}}\ \frac{1}{(4\pi)^{2}}\Big[c\,\phi C_{\mu\nu\rho\sigma}^{2}-a\Big(\phi R_{GB}^{2}-2\hat{G}^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi+(\hat{\square}\phi)^{2}$ (4.28)
@@
$c_{4D,L}=-\frac{1}{30}-c,$
$a_{4D,L}=-\frac{7}{180}-a.$ (4.33)
Here, c-c and a-a are classical contributions coming from (4.31), and the quantum one-loop anomaly from (4.25) is given in (4.22) with $x$ and $y$ in Eq. (4.23):
@@
$\frac{199}{30}+\frac{1}{120}(N_{S}+6N_{F}+12N_{V})-\frac{1}{30}-c=0,$
$\frac{87}{20}+\frac{1}{360}(N_{S}+11N_{F}+62N_{V})-\frac{7}{180}-a=0.$ (4.43)
@@ Appendix C Consistency of the trace anomaly
$\langle T_{\mu}^{\mu}\rangle$ $=$ $\frac{2}{\sqrt{g}}g_{\mu\nu}\frac{\delta S_{\rm eff}}{\delta g_{\mu\nu}}(g_{\mu\nu}=e^{\phi}{\hat{g}}_{\mu\nu})$ (C.1)
$=$ $\int d^{4}x\sqrt{{\hat{g}}}\big(\hat{R}^{2}+A\hat{R}_{\mu\nu}\hat{R}^{\mu\nu}+B\hat{R}_{\mu\nu\lambda\rho}\hat{R}^{\mu\nu\lambda\rho}\big)$
$=$ $\frac{\delta S_{\rm eff}}{\delta\phi},$
@@
$\delta^{2}S_{\rm eff}=\int d^{4}x\sqrt{{\hat{g}}}\Big[-6\hat{R}\,\hat{\square}\delta\phi-A(2\hat{R}^{\mu\nu}\,\hat{\nabla}_{\mu}\hat{\nabla}_{\nu}\delta\phi+\hat{R}\,\hat{\square}\delta\phi)-4B\hat{R}^{\mu\nu}\hat{\nabla}_{\mu}\hat{\nabla}_{\nu}\delta\phi\Big]\wedge\delta\phi.$ (C.3)
@@
$\hat{\nabla}_{\mu}\hat{R}^{\mu\nu}=\frac{1}{2}\hat{\nabla}_{\nu}\hat{R}$, (C.4)
@@
$\delta^{2}S_{\rm eff}=-\int d^{4}\sqrt{{\hat{g}}}\,2(3+A+B)\hat{R}\,\hat{\square}\delta\phi\wedge\delta\phi\Big$. (C.5)

### v2

@@ 2 Conformal anomaly from the conformal gravity
where $\delta\lambda$ is an anticommuting parameter and $c$ is a ghost field for the conformal transformation. [^1] This ghost $c$ and the following antighost $b$ should not be confused with the $a,b,c$ coefficients in Eq. (1.7). The BRST transformation of the ghost is determined by the requirement that the BRST transformation (2.2) be nilpotent. This gives
@@
where anticommuting property of the ghosts is used so that bc=cbbc=-cb. [^2] If we use the transformation (2.4), the same procedure gives $S_{\rm GF+FP}=\int d^{4}x\sqrt{g}\left[-\frac{\alpha}{2}\left(B+cb-\frac{1}{\alpha}R\right)^{2}+\frac{1}{2\alpha}R^{2}+3b\square c\right].$ So practically both transformations give the same result. The original conformal transformation is broken by the gauge fixing term, leaving only the global BRST invariance. Thus we may adopt transformations for the antighost and auxiliary fields different from those suggested by the conformal transformation as long as we consider the proper transformation for the other physical fields. It is interesting that there is no $Rbc$ term in the action (2.9) even though $b$ and $c$ are (fermionic) scalar fields, in contrast to usual bosonic scalar fields (see Appendix B).
@@ 3 Lessons from two-dimensional quantum gravity for conformal gravity – Integrating the trace anomaly
$-\frac{c}{48\pi\epsilon_{2}}\int d^{D}x\sqrt{{\hat{g}}}\,e^{(D-2)\phi/2}\left(\hat{R}-(D-1)\hat{\square}\phi-\frac{(D-1)(D-2)}{4}(\partial_{\mu}\phi)^{2}\right),$ (3.13)
@@ 4.1 Integrating trace anomaly
The Weyl tensor squared in $D$ dimensions is
@@
$\sqrt{g}\,C_{\mu\nu\lambda\sigma}^{2}\to\sqrt{{\hat{g}}}\,e^{((D-4)\phi/2}\hat{C}_{\mu\nu\lambda\rho}^{2}.$ (4.4)
@@
$\frac{c}{(4\pi)^{2}\epsilon}\int d^{D}x\sqrt{g}\,C_{\mu\nu\lambda\sigma}^{2},$ (4.5)
@@
$-\frac{c}{(4\pi)^{2}}\int d^{D}x\sqrt{{\hat{g}}}\,\phi\,\hat{C}_{\mu\nu\lambda\sigma}^{2}.$ (4.6)
@@
$-\frac{a}{(4\pi)^{2}\epsilon}\int d^{D}x\sqrt{g}\,R_{GB}^{2},$ (4.10)
@@
$\frac{a}{(4\pi)^{2}}\int d^{4}x\Big[\phi\,\hat{R}_{\rm GB}^{2}-2\hat{G}^{\mu\nu}\hat{\nabla}_{\mu}\phi\hat{\nabla}_{\nu}\phi+2(\hat{\nabla}^{\mu}\hat{\nabla}^{\nu}\phi)\partial_{\mu}\phi\partial_{\nu}\phi-\frac{1}{4}\{(\partial_{\mu}\phi)^{2}\}^{2}\Big]$
$=\frac{a}{(4\pi)^{2}}\int d^{4}x\sqrt{g}\Big[\phi\,\hat{R}_{\rm GB}^{2}-2\hat{G}^{\mu\nu}\hat{\nabla}_{\mu}\phi\hat{\nabla}_{\nu}\phi-(\partial_{\mu}\phi)^{2}\hat{\square}\phi-\frac{1}{4}\{(\partial_{\mu}\phi)^{2}\}^{2}\Big].$ (4.11)
Collecting, our four-dimensional Liouville action would be
$S_{\rm eff}\hskip-5.69054pt$ $=$ $\hskip-5.69054pt\int d^{4}x\sqrt{{\hat{g}}}\,\Big[\frac{-c}{(4\pi)^{2}}\phi\,\hat{C}_{\mu\nu\lambda\sigma}^{2}+\frac{a}{(4\pi)^{2}}\Big(\phi\,\hat{R}_{\rm GB}^{2}-2\hat{G}^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi-(\hat{\square}\phi)(\partial_{\mu}\phi)^{2}-\frac{1}{4}\{(\partial_{\mu}\phi)^{2}\}^{2}\Big)\Big].$
@@ 4.2 Four-dimensional noncritical conformal gravity
$S_{4D,L}[{\hat{g}},\phi]\hskip-5.69054pt$ $=$ $\hskip-5.69054pt\int d^{4}x\sqrt{{\hat{g}}}\ \frac{1}{(4\pi)^{2}}\Big[-c\phi C_{\mu\nu\rho\sigma}^{2}+a\Big(\phi R_{GB}^{2}-2\hat{G}^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi+(\hat{\square}\phi)^{2}$ (4.28)
@@
$c_{4D,L}=-\frac{1}{30}-c^{\prime},$
$a_{4D,L}=-\frac{7}{180}-a^{\prime}.$ (4.33)
Here, $c^{\prime}$ and $a^{\prime}$ are classical contributions coming from (4.31), and the quantum one-loop anomaly from (4.25) is given in (4.22) with $x$ and $y$ in Eq. (4.23):
@@
$\frac{199}{30}+\frac{1}{120}(N_{S}+6N_{F}+12N_{V})-\frac{1}{30}-c^{\prime}=0,$
$\frac{87}{20}+\frac{1}{360}(N_{S}+11N_{F}+62N_{V})-\frac{7}{180}-a^{\prime}=0.$ (4.43)
@@ Appendix C Consistency of the trace anomaly
$\langle T_{\mu}^{\mu}\rangle$ $=$ $-\frac{2}{\sqrt{g}}g_{\mu\nu}\frac{\delta S_{\rm eff}}{\delta g_{\mu\nu}}(g_{\mu\nu}=e^{\phi}{\hat{g}}_{\mu\nu})$ (C.1)
$=$ $-\frac{\delta S_{\rm eff}}{\delta{\phi}}$
$\sim$ $R^{2}+AR_{\mu\nu}R^{\mu\nu}+BR_{\mu\nu\lambda\rho}R^{\mu\nu\lambda\rho}.$
@@
$\delta^{2}S_{\rm eff}=\int d^{4}x\sqrt{g}\Big[-6R\,\square\delta\phi-A(2R^{\mu\nu}\,\nabla_{\mu}\nabla_{\nu}\delta\phi+R\,\square\delta\phi)-4BR^{\mu\nu}\nabla_{\mu}\nabla_{\nu}\delta\phi\Big]\wedge\delta\phi.$ (C.3)
@@
$\nabla_{\mu}R^{\mu\nu}=\frac{1}{2}\nabla^{\nu}R,$ (C.4)
@@
$\delta^{2}S_{\rm eff}=-\int d^{4}\sqrt{g}\,2(3+A+B)R\,\square\delta\phi\wedge\delta\phi\Big$. (C.5)